8 people shared a prize and receive 125 each. How much would each have receive if there had been only 5 prize winners?
Found 2 solutions by Cromlix, ewatrrr:
Answer by Cromlix(4381)

Total prize = 8 * 125 = 1000
If only 5 prize winners:
1000/5 = 200 each.
Hope this helps.
:-)

Answer by ewatrrr(24785)

 
Hi,
8 people shared a prize and receive 125 each. (8•125 = 1000 Total)
1000/5 = 200 each if there had been only 5 prize winners